OpenStudy (anonymous):

In a two-digit number, the sum of the digits is 14. Twice the tens digit exceeds the units digit by one. Find the number.

OpenStudy (anonymous):

59

OpenStudy (anonymous):

say the no is 10y+x now as per condition x+y=14.........(i) 2y=x+1......(ii) so the ans is 59..

OpenStudy (anonymous):

59